OpenStudy (yourboydusty):

What is the slope of the line given by the equation below? y-5=3/5(x-10)

OpenStudy (yourboydusty):

A. -10 B. 6 C. 10 D. - E. F. -6

OpenStudy (yourboydusty):

D.-5/3 E.3/5

OpenStudy (yourboydusty):

@3mar

OpenStudy (3mar):

Can you simplify it a little bit?

OpenStudy (yourboydusty):

what do you mean?

OpenStudy (3mar):

Ok like that

OpenStudy (3mar):

\[y-5=\frac{ 3 }{ 5 }(x-10)\]

OpenStudy (3mar):

and then \[y-5=\frac{ 3 }{ 5 }x-6\]

OpenStudy (3mar):

and then \[y=\frac{ 3 }{ 5 }x-1\]

OpenStudy (3mar):

What do you think?

OpenStudy (yourboydusty):

i think the answer is going to be E.

OpenStudy (3mar):

which is?

OpenStudy (yourboydusty):

3/5

OpenStudy (yourboydusty):

???

OpenStudy (3mar):

GREAT

OpenStudy (3mar):

Does E say that?

OpenStudy (yourboydusty):

yes

OpenStudy (3mar):

so it is

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (3mar):

Any more questions?